# Weight-Space Linear Recurrent Neural Networks

### v1

@@ 2.3 Training & Inference
Analogous to SSMs Gu et al. (2021) and subsequent linear recurrence architectures Orvieto et al. (2023); Movahedi et al. (2025), WARP supports dual training modes: convolutional and recurrent. The former is accomplished through a systematic unrolling of the linear recurrence formulated in Equation 1, enabling the derivation of a convolution kernel $K$ such that $\theta_{0:T}=K\star\Delta\mathbf{x}_{0:T}$. Comprehensive notational specifications, algorithms, and rigorous mathematical derivations are elaborated in Appendix A. In recurrent mode, we distinguish the auto-regressive (AR) and the relatively memory-expensive [^3] Although equal to AR in computational complexity, the recurrent non-AR setting requires more memory because like the convolutional mode, it materializes all high-dimensional hidden states $\theta_{t}$. non-AR settings. The non-AR setting never sees its own predictions, making it ideal for classification tasks wherein $\theta_{t}(\cdot)$ only generates logits.
The recurrent AR setting is particularly advantageous for noisy forecasting tasks that necessitate accurate modeling of the sequential data distribution $p(\mathbf{y}_{t}|\mathbf{y}_{<t})$. To mitigate exposure bias (Schmidt, 2019), we implement teacher forcing with scheduled sampling Bengio et al. (2015), wherein the model incorporates uncertainties by sampling $\hat{\mathbf{y}}_{t}\sim\mathcal{N}(\boldsymbol{\hat{\mu}}_{t},\boldsymbol{\hat{\sigma}^{2}_{t}})$ using the reparametrization trick [^4] We remark that this sampling is not required during inference on smooth sequences. Kingma et al. (2013). Selection between ground truth $\mathbf{y}_{t}$ and predicted $\hat{\mathbf{y}}_{t}$ follows a Bernoulli distribution with probability $p_{\text{forcing}}$, which we define as a training hyperparameter. That said, we consistently use $\hat{\mathbf{y}}_{t-1}$ in the input difference seen in (1).
During inference on regression problems, the model operates fully auto-regressively, i.e. $p_{\text{forcing}}=1$ within the context window, and $p_{\text{forcing}}=0$ in the forecast window, regardless of the training mode that was used. Although other functions can be considered, our WARP models are trained by minimizing either the mean-squared error (MSE) for deterministic predictions, or the simplified negative log-likelihood (NLL) for probabilistic predictions:
$\mathcal{L}_{\text{MSE}}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|_{2}^{2},\qquad\mathcal{L}_{\text{NLL}}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}\left(\frac{\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|_{2}^{2}}{2\hat{\boldsymbol{\sigma}}_{t}^{2}}+\log\hat{\boldsymbol{\sigma}}_{t}\right).$ (3)
@@ 3 Experiments
MNIST
$L=\mathbf{100}$ $L=\mathbf{300}$ $L=\mathbf{600}$
@@
CelebA
$L=\mathbf{100}$ $L=\mathbf{300}$ $L=\mathbf{600}$
@@
We evaluate WARP on real-world multivariate time series datasets, 2D images, and physical systems. Our experiments elucidate empirical questions regarding forecasting, classification, and dynamical system reconstruction and generalization. Experimental details can be found in Appendix C.
3.1 Image Completion & Time Series Forecasting
In the first part of our experiments, we focus on forecasting applied first to raster-scanned pixel-by-pixel image completion, followed by real-world electricity prediction.
@@ 3.1 Image Completion & Time Series Forecasting
Image completion is cast as a prediction of pixel intensities. We focus on two datasets: MNIST handwritten digits LeCun et al. (1998), and the celebrity face attributes CelebA Liu et al. (2015)—additional image datasets are considered in Appendix D. 2D images are flattened into 1D sequences with length $T=784$ for MNIST and $T=1024$ for CelebA.
Following Rush and Karamcheti (2022), the completion task is conditioned on contexts of variable length $L$. We compare WARP against long-established baselines like GRU Cho et al. (2014) and LSTM Hochreiter and Schmidhuber (1997); against state-of-the-art (SoTA) SSMs like S4 Gu et al. (2021); and against the ConvCNP convolution-based meta-learning baseline Gordon et al. (2020) specifically designed for image completion. All models are trained with the NLL loss in recurrent AR mode to ensure fair comparison [^5] Additionally, the traditional dual RNN encoder-decoder approach from Cho et al. (2014) is not used with GRU or LSTM due to the need for variable $L$ during testing.; and feature nearly the same number of learnable parameters: approximately 1.68M for MNIST, and 2M for CelebA. Results in Table 1 demonstrate the generative performance of WARP as measured by the MSE and the uncertainty-aware bits-per-dimension (BPD) metrics. We focus on the top performing models across three runs, with corresponding qualitative results — best captured by the BPD — in Appendix E.
Figure 2: Heatmap of test MSEs on the ETT task.
@@
We evaluate WARP’s performance on long-range energy forecasting tasks with the Electricity Transformer Temperature (ETT) dataset Zhou et al. (2021). Following established methodological protocols Nanbo et al. (2025), we utilize the open-source TSLib [^6] TSLib is accessible at https://github.com/thuml/Time-Series-Library.git to obtain preprocessed data splits which we further normalize using train set statistics (additional data processing details can be found in Appendix B).
The models are tasked with predicting 96 time steps based on a context of length $L=96$, with performance evaluated using the mean MSE across three runs. The results are shown in Figure 2, where the best along the columns is reported in a box while the second-best is underlined. It demonstrates WARP’s superiority, achieving the best performance on all subsets except the ETTh1 subset where it ranked second. These results are particularly noteworthy given WARP’s straightforward design. Indeed, WARP offers an elegant balance between architectural simplicity and predictive power. Additional results on the ETT dataset are presented in Appendix D.
@@ 3.2 Dynamical System Reconstruction
As our final forecasting benchmark, we evaluate WARP’s capabilities on dynamical system reconstruction (DSR) Göring et al. (2024). The experiments presented in this section highlight the challenge of OoD generalization to physical parameters, a research area that has recently experienced a significant surge in interest Nzoyem et al. (2025b); Brenner et al. (2025).
We establish four DSR benchmark datasets: $\bullet$ (1) Mass Spring Damper (MSD) characterizes challenging damped oscillatory dynamics through physical parameters $(m,k,c)$, with trajectories of length $T=256$, of which $L=100$ serve as context; $\bullet$ (2) MSD-Zero is a version of MSD which varies, in addition to the significant relative scales and wide ranges of $(m,k,c)$, the initial condition $\mathbf{x}_{0}$; $\bullet$ (3) Lotka-Volterra (LV) is parametrized by coefficients $(\alpha,\beta,\gamma,\delta)$; $\bullet$ (4) SINE comprises sine curves $\tau\mapsto\sin(2\pi\tau+\varphi)$ with varying phases $\varphi$ (we set $T=16$ and $L=1$, resulting in an initial value problem). Each test set incorporates out-of-distribution parameters, except for SINE, which primarily tests model performance under sample size constraints. Comprehensive data generation protocols for all four datasets are detailed in Appendix B. We benchmark against two established RNNs and the Time Series Transformer (TST) from HuggingFace Jain (2022) specialized for forecasting. We evaluate WARP in a black-box setting — which embeds no explicit physical knowledge in the root network — followed by the more interpretable grey-box.
Table 2: Average test MSE ($\downarrow$) and MAE ($\downarrow$) for dynamical system reconstruction. Best results are reported in bold, while the second-best are underlined. All are reported with ×$\times 10^{-2}$ scaling, except for SINE* with ×$\times 10^{-4}$. SINE* indicates that metrics are computed upon training on its “Small” data split. WARP-Phys indicates the variant of WARP with physical constraints in the root network.
@@
Our results, presented in Table 2, highlight how weight-space linear RNNs consistently outperforms all baseline models across problem domains. Importantly, the standard WARP configuration which uses a root black-box MLP ranks within the top two in three out of four problem settings. We observe that TST — denoted simply as Transformer — exhibits significant performance degradation on the SINE* dataset (which comprises only 10 sequences), corroborating the documented limitation of Transformer models to overfit in data-scarce regimes due to their inherently high parameter complexity Dosovitskiy et al. (2020).
@@
A principal advantage of WARP is its capacity to incorporate domain-specific knowledge into the root network, exemplified on the SINE* experiment by embedding the explicit mathematical formulation $\tau\mapsto\sin(2\pi\tau+\hat{\varphi})$ in its forward pass, where $\hat{\varphi}$ is predicted by a MLP. The resulting architecture, WARP-Phys, demonstrates substantial performance improvements relative to WARP (more than one order of magnitude on MSD). Notably, the incorporation of such a powerful physical prior on SINE* underscores the value of an expressive but data-efficient initial network $\phi$ whose task it is to capture a representation of $\varphi$. Indeed, all models including WARP and WARP-Phys perform poorly on the extreme “Tiny” data split (not reported in Table 2). We provide additional details as ablations in Sections 3.4 and D.
@@
We evaluate our model’s pattern memorization capabilities on the Lotka-Volterra (LV) dataset, which constitutes a continuous analogue of the established repeat-copy benchmark Tay et al. (2021); Orvieto et al. (2023). To generate the output shown in red in LABEL:fig:lotka_inout, we triplicate a concise segment of the input, delineating them by a 10-token long sequence of $-1$s. In this challenging problem, WARP demonstrates superior performance relative to all baselines, with the GRU achieving the second-highest performance metrics (see Table 2). These findings suggest that the high-resolution weight-space state representation exhibits enhanced pattern retention capabilities compared to conventional methodologies. We note that this particular evaluation protocol is incompatible with the WARP-Phys variant due to the deliberate introduction of artificial discontinuities in the temporal sequences. Comprehensive analyses of additional results pertaining to this task, alongside other dynamical system reconstruction benchmarks, are presented in Appendix D.
@@ 3.3 Multivariate Time Series Classification
We now consider the classification setting. We consider six datasets from the University of East Anglia (UEA) multivariate time series classification archive (UEA-MTSCA) Bagnall et al. (2018). The six datasets are selected and preprocessed following the criteria of known difficulty for deep sequence models and data abundance, with sequence length ranging from 405 to almost 18k Walker et al. (2024). Our model is compared to both discrete and continuous recurrent baselines Morrill et al. (2021); Kidger et al. (2020); Orvieto et al. (2023); Smith et al. (2023); Walker et al. (2024); Gu and Dao (2024). All models are trained, validated and tested with the 70:15:15 split. Additional details on the dataset preprocessing and the baselines is provided in Appendix B.
Table 3 presents test accuracy metrics across all six benchmark datasets for WARP (trained in recurrent non-AR mode) and competing methodologies as reported in Walker et al. (2024). Our analysis reveals that WARP demonstrates exceptional performance across the majority of tasks, establishing new state-of-the-art benchmarks on the SCP2 dataset, and competitive second-place on two datasets: Ethanol and Heartbeat. Despite these strong results, we observe that our framework encounters performance limitations when processing extremely long sequences (specifically EigenWorms and Motor), potentially attributable to the well-documented vanishing gradient phenomenon that characteristically affects recurrent architectures Zucchet and Orvieto (2024). These empirical findings substantiate WARP’s efficacy as a robust classification framework for diverse real-world time series applications.
@@
WARP 60.98 $\pm$ 3.1 80.00 $\pm$ 2.0 57.89 $\pm$ 1.5 31.65 $\pm$ 0.8 77.42 $\pm$ 2.2 50.88 $\pm$ 2.3
3.4 Ablation Studies
We briefly discuss several experiments carried out to gain insights into our model. For all ablation studies, experimental protocols like training hyperparameters are presented in Appendix C, with detailed quantitative results in Appendix D.
Eliminating the root network
The root network $\theta_{t}$ is integral to the efficacy of WARP. Although not an absolute prerequisite for the WARP-Phys variant, it nonetheless persists as a pivotal constituent of our framework. Illustratively, the omission of $\theta_{t}$ in favor of directly fitting $\varphi$ for the SINE modeling problem results in a catastrophic degradation in model expressivity.
Root network evaluation
When we fix the evaluation point $\tau$ throughout training, we observe mild degradation in the qualitative results. This emphasizes the importance of the diagonal decoding “readout” direction $\theta_{t}(\tau)$, illustrated for the MSD problem in LABEL:fig:readout_matrix.
Initial network configuration
Since WARP’s weight trajectories is driven by the changes in the signal and not the signal itself (see Equation 1), it is important to have an expressive initial hypernetwork $\phi:\mathbf{x}_{0}\mapsto\theta_{0}$, which embeds the initial tokens into suitable weight-spaces Kidger (2022). Our empirical investigations reveal that sidestepping this component substantially curtails the model performance on complex synthetic benchmarks, such as MSD-Zero, and on real-world datasets, including ETT.
Figure 5: Data efficiency comparison on the SINE datasets.
Data efficiency
With $L=1$, the SINE benchmark is a challenging initial value problem. At equal (root) neural network parameter counts, we vary the number of training samples, and we plot MSE and MAPE test metrics for WARP and the Neural ODE Chen et al. (2018). The results, depicted in Figure 5, not only show improved performance across data-regimes, but they also indicate that more data is not necessarily better for WARP’s performance, suggesting potential for monotone learning Bousquet et al. (2022).
Dense state transitions & Channel mixing
4 Related Work

### v2

@@ 2.3 Training & Inference
Analogous to SSMs [36] and subsequent linear recurrence architectures [75; 69], WARP supports dual training modes: convolutional and recurrent. The former is accomplished through a systematic unrolling of the linear recurrence formulated in Eq. 1, enabling the derivation of a convolution kernel $K$ such that $\theta_{0:T}=K\star\Delta\mathbf{x}_{0:T}$. Comprehensive notations, algorithms, and rigorous mathematical derivations are elaborated in Section B.2.2. In recurrent mode, we distinguish the auto-regressive (AR) and the relatively memory-expensive [^3] Although equal to AR in computational complexity, the recurrent non-AR setting requires more memory because, like the convolutional mode, it materialises all high-dimensional hidden states $\theta_{t}$. non-AR settings. The non-AR setting never sees its own predictions, making it ideal for classification tasks wherein $\theta_{t}(\cdot)$ only generates logits.
The recurrent AR setting is particularly advantageous for noisy forecasting tasks that necessitate accurate modelling of the sequential data distribution $p(\mathbf{y}_{t}|\mathbf{y}_{<t})$. To mitigate exposure bias [81], we implement teacher forcing with scheduled sampling [11], wherein the model incorporates uncertainties by sampling $\hat{\mathbf{y}}_{t}\sim\mathcal{N}(\bm{\hat{\mu}}_{t},\bm{\hat{\sigma}^{2}_{t}})$ using the reparametrisation trick [^4] We remark that this sampling is not required during inference on smooth sequences like dynamical systems. [54]. Selection between ground truth $\mathbf{y}_{t}$ and predicted $\hat{\mathbf{y}}_{t}$ follows a Bernoulli distribution with probability $p_{\text{forcing}}$, which we define as a training hyperparameter. That said, we consistently use $\hat{\mathbf{y}}_{t-1}$ in the input difference seen in Eq. 1.
During inference on regression problems, the model operates fully auto-regressively, i.e., $p_{\text{forcing}}=1$ within the context window, and $p_{\text{forcing}}=0$ in the forecast window, regardless of the training mode.
Although other loss functions can be considered, our WARP models are trained by minimizing either the mean-squared error (MSE) for deterministic predictions, or the simplified negative log-likelihood (NLL) for probabilistic predictions:
$\mathcal{L}_{\text{MSE}}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|_{2}^{2},\qquad\mathcal{L}_{\text{NLL}}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}\left(\frac{\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|_{2}^{2}}{2\hat{\bm{\sigma}}_{t}^{2}}+\log\hat{\bm{\sigma}}_{t}\right).$ (2)
@@
Our learning pseudocodes are detailed in Algorithms 1 and 2 of Appendix B, outlining the strong connection to the fast weights and test-time training literatures [80; 7; 98]. At each training step, the slow-changing RNN parameters $A,B$ and $\phi$ (or $\theta_{0}$) are updated once using gradient descent to minimise one of the loss objectives above. The fast-changing weights $\theta_{t}$, however, are updated T1T-1 times using Eq. 1, i.e., not using gradient descent. This distinction is central to our model’s gradient-free adaptation process.
@@ 3 Experiments
We evaluate WARP on real-world multivariate time series datasets, 2D images, and physical systems. Our experiments elucidate empirical questions regarding forecasting, classification, and dynamical system reconstruction and generalisation. Additional experiments allow us to demonstrate WARP’s in-context learning abilities. Theoretical results are presented in Section B.2, and experimental details can be found in Appendix D.
3.1 Image Completion, Energy Prediction & Traffic Forecasting
@@
MNIST $L=\mathbf{100}$ $L=\mathbf{300}$ $L=\mathbf{600}$
@@
CelebA $L=\mathbf{100}$ $L=\mathbf{300}$ $L=\mathbf{600}$
@@
In the first part of our experiments, we focus on forecasting applied first to raster-scanned pixel-by-pixel image completion, followed by real-world electricity and traffic flow.
@@
Image completion is cast as a prediction of pixel intensities. We focus on two datasets: MNIST handwritten digits [58], and the celebrity face attributes CelebA [61]—additional image datasets are considered in Appendix E. 2D images are flattened into 1D sequences with length $T=784$ for MNIST and $T=1024$ for CelebA.
Following [79], the completion task is conditioned on contexts of variable length $L$. We compare WARP against long-established baselines like GRU [18] and LSTM [43]; against state-of-the-art (SoTA) SSMs like S4 [36]; and against the ConvCNP convolution-based meta-learning baseline [30] specifically designed for image completion. All models are trained with the NLL loss in recurrent AR mode to ensure fair comparison, and feature nearly the same number of learnable parameters: approximately 1.68M for MNIST, and 2M for CelebA. Results in Table 1 demonstrate the generative performance of WARP as measured by the MSE and the uncertainty-aware bits-per-dimension (BPD) metrics. We focus on the top performing models across three runs, with corresponding qualitative comparisons — best captured by the BPD — in Appendix F. For instance, LABEL:fig:mnist_comp_vis shows that at small parameter count, WARP is the only model to accurately generate digits without substantial artefacts.
(a) (a) Comparison of a GRU [18], LSTM [43], S4 [79], and WARP on the MNIST image completion task with $L=300$ initial pixels. All models are roughly at the same size of 1.7M parameters, with architectures described in Section C.2. The leftmost column represents target images with context (in white) and ground truths (in green). Predicted forecasts are drawn in red. (b) Heatmap of test MSEs ($\downarrow$) on the ETT task, with best results enclosed and second-best underlined.
@@
We evaluate WARP’s performance on long-range energy forecasting tasks with the Electricity Transformer Temperature (ETT) dataset [105]. Following established methodological protocols [70], we utilise the open-source TSLib [^5] https://github.com/thuml/Time-Series-Library.git to obtain preprocessed data splits which we further normalise using train set statistics (additional data processing details can be found in Appendix C).
The models are tasked with predicting 96 time steps based on a context of length $L=96$, with performance evaluated using the mean MSE across three runs. The results are shown in LABEL:fig:ettresults, where the best along the columns is reported in a box while the second-best is underlined. It demonstrates WARP’s superiority, achieving the best performance on all subsets except the ETTh1 subset, where it ranked second. These results are particularly noteworthy given WARP’s straightforward design. Indeed, WARP offers an elegant balance between architectural simplicity and predictive power. Additional results on the ETT dataset are presented in Appendix E.
Table 2: Performance on PEMS08 [85]. SoTA baselines leverage spatial information, as reported in [60].
MODEL MAE RMSE
GMAN [100] 14.57 24.71
D2STGNN [101] 14.35 24.18
STIDGCN [60] 13.45 23.28
WARP 6.59 10.10
Traffic Flow Forecasting
We conduct extensive experiments on the PEMS08 real-world traffic network [85]. The network consists of 170 nodes, from which 3 features are collected at 5-minute intervals over two months. The standard task is to predict the traffic flow for the next hour (12 steps) given the historical data from the previous hour (12 steps). Given its chunk-wise forecasting — which significantly differs from the setting in Fig. 2 — we employ the non-AR mode to train and test WARP. Additionally, we preprocess the input sequence with a non-causal convolution, as detailed in Appendix D.
As demonstrated in Table 2, our model achieves a MAE of 6.59 and a RMSE of 10.10. These results represent a significant improvement over the current state-of-the-art on the PEMS08 benchmark [60], reducing MAE by over 50% compared to the best-published model. It is particularly noteworthy that our model achieves this performance without using the inherent graph structure, outperforming complex Attention and Graph Neural Network (GNN) architectures that are specifically designed to leverage this spatial information.
@@ 3.2 Dynamical System Reconstruction
As our final forecasting benchmark, we evaluate WARP’s capabilities on dynamical system reconstruction (DSR) [31]. The experiments presented in this section highlight the challenge of OoD generalisation to physical parameters, a research area that has recently experienced a significant surge in interest [74; 15].
We establish four DSR benchmark datasets: $\bullet$ (1) Mass Spring Damper (MSD) characterises challenging damped oscillatory dynamics through physical parameters $(m,k,c)$, with trajectories of length $T=256$, of which $L=100$ states serve as context; $\bullet$ (2) MSD-Zero is a version of MSD which varies, in addition to the significant relative scales and wide ranges of $(m,k,c)$, the initial condition $\mathbf{x}_{0}$; $\bullet$ (3) Lotka-Volterra (LV) is parametrised by coefficients $(\alpha,\beta,\gamma,\delta)$; $\bullet$ (4) SINE comprises sine curves $\tau\mapsto\sin(2\pi\tau+\varphi)$ with varying phases $\varphi$ (we set $T=16$ and $L=1$, resulting in an initial value problem). Each test set incorporates out-of-distribution parameters, except for SINE, which primarily tests model performance under sample size constraints. Comprehensive data generation protocols for all four datasets are detailed in Appendix C. We benchmark against two established RNNs and the Time Series Transformer (TST) from HuggingFace [45] specialised for forecasting. We evaluate WARP in a black-box setting — which embeds no explicit physical knowledge in the root network — followed by the more interpretable grey-box.
Table 3: Average test MSE ($\downarrow$) and MAE ($\downarrow$) for dynamical system reconstruction. Best results are reported in bold, while the second-best are underlined. All are reported with $\times 10^{-2}$ scaling, except for SINE* with $\times 10^{-4}$. SINE* indicates that metrics are computed upon training on its “Small” data split. WARP-Phys indicates the variant of WARP with physical constraints in the root network.
@@
Our results, presented in Table 3, highlight how weight-space linear RNNs consistently outperform all baseline models across problem domains. Importantly, the standard WARP configuration, which uses a root black-box MLP, ranks within the top two in three out of four problem settings. We observe that TST — denoted simply as Transformer — exhibits significant performance degradation on the SINE* dataset (which comprises only 10 sequences), corroborating the documented limitation of Transformer models to overfit in data-scarce regimes due to their inherently high parameter complexity [26].
@@
[…]
Figure 4: Sample LV input/output.
@@
We evaluate our model’s pattern memorisation capabilities on the Lotka-Volterra (LV) dataset, which constitutes a continuous analogue of the established repeat-copy benchmark [86; 75]. To generate the output shown in red in Fig. 4, we triplicate a concise segment of the input, separating the repetitions by a 10-token long sequence of 1-1s. In this challenging problem, WARP demonstrates superior performance relative to all baselines, with the GRU achieving the second-highest performance metrics (see Table 3). These findings suggest that the high-resolution weight-space state representation exhibits enhanced pattern retention capabilities compared to conventional methodologies. We note that this particular evaluation protocol is incompatible with the WARP-Phys variant due to the deliberate introduction of artificial discontinuities in the temporal sequences. Comprehensive analyses of additional results pertaining to this task, alongside other dynamical system reconstruction benchmarks, are presented in Appendix E.
@@ 3.3 Multivariate Time Series Classification
We now consider the classification setting. We consider six datasets from the University of East Anglia (UEA) multivariate time series classification archive (UEA-MTSCA) [8]. The six datasets are selected and preprocessed following the criteria of known difficulty for deep sequence models and data abundance, with sequence length ranging from 405 to almost 18k [93]. Our model is compared to both discrete and continuous recurrent baselines [68; 53; 75; 84; 93; 35]. All models are trained, validated, and tested with the 70:15:15 split. Additional details on the dataset preprocessing, the baselines, and the positional encoding used for $\tau$ are provided in Appendix C.
Table 4 presents test accuracy metrics across all six benchmark datasets for WARP (trained in recurrent non-AR mode) and competing methodologies as reported in [93]. Our analysis reveals that WARP demonstrates exceptional performance across the majority of tasks, establishing new state-of-the-art accuracies on the SCP2, Ethanol, and Heartbeat datasets, and competitive top three on 5 datasets. Despite not being designed with long-range dependencies in mind, WARP displays impressive potential on extremely long sequences such as EigenWorms and Motor, outperforming established models such as Mamba [35] and NCDE [53]. This overcoming of the well-documented vanishing and exploding gradient problems in recurrent architectures [109] is attributed to our careful initialisation scheme in Section 2.2, and the positional encoding scheme using sines and cosines with variable frequencies [89]. These empirical findings substantiate WARP’s efficacy as a robust classification framework for diverse real-world time series applications.
@@
WARP 70.93 $\pm$ 2.7 83.53 $\pm$ 2.0 57.89 $\pm$ 1.4 36.49 $\pm$ 2.8 80.65 $\pm$ 1.9 56.14 $\pm$ 5.1
3.4 In-Context Learning with Randomly Generated Keys
A key strength of WARP is illustrated in the classical in-context learning (ICL) setting of [99], where the objective is to learn a linear mapping from $N$ randomly generated keys $\mathbf{x}_{i}\in\mathbb{R}^{D_{x}-1}$ and their corresponding values $y_{i}\in\mathbb{R}^{1}$. In this setup, WARP aims to learn the weights of the root network that approximate this mapping. We adapt the task by transforming the input sequence into its cumulative sum along the time dimension, and predicting the value corresponding to the query $\mathbf{x}_{q}$ (see LABEL:fig:icl_transform_sub). This preserves the underlying function while allowing the model to exploit key-value pairs dependencies. WARP is trained in its recurrent, non-autoregressive mode with a MSE loss over the entire 1D output sequence of length $T=N+1=32$. The results, shown in LABEL:fig:allvalues and LABEL:fig:queryonly, highlight WARP’s ability to perform sub-quadratic in-context learning and generalize effectively.
(a) Pipeline and results for in-context learning. (a) Cumulative sum transformation and subsequent processing of the input matrix. (b) Linear mappings learned between scalar keys and values of the same sequences ($D_{x}=2$). (c) Ground truth vs. query point predictions ($D_{x}=8$).
A key advantage of this approach is that once the model has learned from the context, the final root network $\theta_{T-1}:\sum_{i=1}^{N}\mathbf{x}_{i}+\mathbf{x}_{q}\mapsto\sum_{i=1}^{N}\hat{y}_{i}+\hat{y}_{q}$, which is equivalent to $\theta_{T-1}:\mathbf{x}_{q}\mapsto\hat{y}_{q}$, can be extracted. This allows it to process subsequent queries without needing to re-evaluate the entire sequence from scratch. This method yields significant computational savings compared to other models capable of ICL [59].
4 Discussion & Conclusion
4.1 Core Advantages
WARP demonstrates outstanding results across a multitude of data modalities, both in-distribution and out-of-distribution, as evidenced by the extensive empirical results on time series forecasting and classification we have presented (see Tables 1, LABEL:fig:mnist_comp_vis, LABEL:fig:ettresults, 2, 3, 4 and 7(a)). Additional results showcasing a 93% classification accuracy on sequential MNIST, along with ablation studies and further results on synthetic datasets are provided in Appendix E. Specifically, Section E.3 illustrates the excellent computational efficiency of our approach, as measured by wall-clock training time per epoch, peak GPU usage, and parameter counts.
By letting the data directly interact with the weights as in Eq. 1, WARP showcases the appealing in-context learning ability to fine-tune an auxiliary network without gradients at test-time [10; 94]. Additionally, WARP is the latest scientific machine learning [20] technique that seamlessly integrates interpretable physical knowledge into its predictions, a feature standard RNNs have overlooked. This demonstratively allows for sample-efficient training and improved generalisation. Finally, the WARP architecture, through its input difference, bears resemblance to synaptic plasticity in biological neural networks [16]. This neuromorphic quality enables more biologically plausible learning dynamics.
4.2 Limitations
Some design decisions that strengthen WARP equally face limitations that we outline as promising avenues for future work. First, the size of the matrix $A$ limits scaling to huge root neural networks. Our experiments conducted on a RTX 4080 GPU with 16GB memory could only support moderate $D_{\theta}$ values, leaving open the question of how expressive WARP models can become if scaled further. Second, more theoretical research is needed to supplement the current state of the weight-space learning literature. Our work remains mostly empirical, despite introducing theory-informed algorithms in Section B.2 and leveraging the underpinnings of NCDEs as universal approximators generalizing RNNs in continuous time settings [51]. Future work would seek such first principles to reduce the memory footprint of the matrix $A$, which could simultaneously enable low-rank or diagonal parametrisation [37] or neuron permutation equivariance [102].
4.3 Conclusion
In this work, we introduced Weight-Space linear RNNs, a novel family of sequence models that operates directly within the weight space of neural networks, offering a distinct paradigm from traditional recurrent architectures. We argue that the high-dimensional weight space can be used for intermediate representations, resulting in infinite-dimensional RNN hidden states and high-capacity memory. Our comprehensive experiments demonstrate that our models exhibit superior expressivity and generalisation capabilities, enabling a powerful form of gradient-free adaptation in response to sequential input differences, and showing exceptional abilities when integrating domain-specific knowledge from physical systems. Our framework draws intriguing parallels to neuromorphic learning principles, leading us a step further towards human-level artificial intelligence.
Broader Impact
While their benefits are evident from Section 4.1, malicious deployment of our self-adaptable models in scenarios they were not designed for could lead to serious adverse outcomes. Additionally, high-energy costs from high-dimensional weight-space computations could increase disparities in our field. To limit the potential for such outcomes and to improve the democratisation of AI, our data and models are openly available at https://anonymous.4open.science/r/warp.
Acknowledgments
This work was supported by UK Research and Innovation grant EP/S022937/1: Interactive Artificial Intelligence, and EPSRC program grant EP/R006768/1: Digital twins for improved dynamic design. We thank Hengshuai Yao and Yasin Abbasi-Yadkori for valuable discussions culminating in ideas that helped improve the appeal and performance of weigh-space linear recurrent neural networks.
Appendix A Related Work